# Multi-Crit: Benchmarking Multimodal Judges on Pluralistic Criteria-Following

### v1

@@ Abstract
**footnotetext: Core contributors.
@@ 1 Introduction
The recent advancements of large multimodal models (LMMs) [li2024llava, liu2023visual, bai2025qwen2, chen2024internvl, chen2025eagle, hurst2024gpt, an2025llava] in visual understanding [tong2024cambrian, yu2016modeling, yu2023mm, guan2024hallusionbench] and reasoning [meng2025mm, zhang2024mathverse, jiang2025mme] have underscored the growing need for scalable and reliable multimodal evaluation systems. However, evaluating free-form outputs in open-ended or reasoning-intensive tasks remains inherently challenging, as static or reference-based metrics often fail to generalize and provide limited insight into model behavior. A growing trend has emerged to deploy LMMs themselves as generative judges (LMM-as-a-Judge[zhang2023gpt, xiong2025llava]. Given a multimodal prompt consisting of an image and a user query, model responses, and a set of predefined evaluation criteria, the judge model produces either scalar scores for individual responses or pairwise preference judgments, accompanied by textual justifications for its evaluation. Owing to its scalability and flexibility in defining task-specific criteria, this paradigm has been widely adopted for open-ended evaluation across diverse multimodal benchmarks [lu2024wildvision, guan2024hallusionbench, zhang2024mathverse, ye2025painting, yu2025rlaif, li2024llava, jiang2025mme, sun2023aligning, xie2025vlms]. A series of works fine-tune open-source models to develop dedicated judge and critic capacities, and have proven effective in providing AI feedback for aligning LMMs through reinforcement fine-tuning [li2024vlfeedback, wang2025unified, li2024vlfeedback, liu2025improving, wang2025enhancing] or test-time scaling [zhang2025r1, wang2025llava].
The reliability of the LMM-as-a-Judge paradigm relies on two key components: (1) consistency with human judgment, and (2) flexibility to follow diverse, task-specific evaluation criteria that capture different aspects of model behavior (e.g., factuality, visual grounding, structural coherence). However, existing multimodal reward and judge benchmarks [li2025vl, zhang2025mm, chen2024mllm, zeng-etal-2025-mm] mainly address the first aspect by providing a single overall preference label for each pair of model responses. This coarse formulation overlooks the evaluation of judge models in producing criterion-specific judgment. Specifically, two responses may exhibit trade-offs across different criteria—for instance, a concise but factually correct answer versus a visually grounded yet slightly hallucinated one. This motivates a key question: Can LMM-based judges follow diverse evaluation criteria and recognize conflicts among criterion-level judgments?
@@
    How to construct evaluation data with multi-criterion human preferences that capture diverse aspects of response quality and reveal intra-sample conflicts? We design a rigorous data curation and preference-annotation pipeline that: (1) collects diverse multimodal prompts spanning both open-ended content generation and verifiable reasoning tasks; (2) selects challenging response pairs from various strong LMMs with fine-grained criterion-level quality differences via multi-step filtering; (3) defines clear evaluation criteria covering complementary aspects of multimodal judgment; and (4) gathers high-quality human annotations ensuring reliability and inter-annotator consistency.
@@
Overall, Multi-Crit serves as the first benchmark for assessing multimodal judges’ adherence to pluralistic evaluation criteria, built upon a curated dataset, novel metrics, and analyses that uncover key limitations and directions for more reliable, steerable evaluation systems.
@@ 2 Related Work
Large Multimodal Model as Judges and Critics.
Large multimodal models (LMMs) are widely adopted for evaluating model-generated responses, providing assessments ranging from scalar [li2024llava, yu2023mm, zhang2024mathverse, guan2024hallusionbench] to pairwise preference [liu2023visual, lu2024wildvision, yu2025rlaif]. GPT-4V [zhang2023gpt] was first shown to align closely with human judges across diverse multimodal evaluation tasks, motivating efforts to develop open-source counterparts. Subsequent works fine-tuned open-source LMMs into either BT-style reward models with dedicated reward heads [sun2023aligning, zang2025internlm, zhang2025basereward] or generative judges that produce textual justifications and scores [lee2024prometheus, xiong2025llava, zhang2025mm, zhang2025r1, wang2025unified, wang2025llava] in an autoregressive manner. While our benchmark centers on generative judge models—due to their flexible instruction-following abilities inherent in the autoregressive generation process of LMMs—the proposed methodology is broadly applicable across diverse reward modeling paradigms.
Benchmarking Multimodal Judges.
Benchmarking reward and judge models is crucial for evaluating the reliability of AI feedback, which underpins both automated evaluation and the enhancement of LMM performance through preference alignment. MLLM-as-a-Judge [chen2024mllm] is the first to evaluate LMM as judges on academic tasks; VL-Rewardbench [li2025vl], multimodal-rewardbench [yasunaga2025multimodal] and MM-RLHF [zhang2025mm] Benchs extend prompts sources to diverse real-world scenarios, reasoning and safety tasks. Following reward and judge benchmarks in text-only LLMs [lambert2024rewardbench, tan2024judgebench, zhou2024rmb], these multimodal benchmarks evaluate judges by their overall preference between paired responses, measured against human or verifier-provided golden labels. MM-Critic further extends evaluation to assess the correctness of generated textual critiques by leveraging GPT-generated scoring rubrics [zeng-etal-2025-mm]. However, none of these benchmarks evaluate multimodal judges on their ability to follow diverse criteria for each response pair, and fine-grained, criterion-level evaluation remains underexplored.
Criteria-Following Judgment.
When deploying model-based judges [zhang2024mathverse, yu2023mm, liu2023visual, yu2025rlaif, xiong2025llava], existing benchmarks typically define task-specific criteria, such as helpfulness [lu2024wildvision], visual grounding [ye2025painting], or hallucination detection [sun2023aligning]. While recent work has begun evaluating criterion-following in the textual domain—either by embedding criterion-level differences into responses [pitis2024improving] or summarizing from human justifications [wang2025rlbff]—we extend this line of research to the multimodal setting, providing a comprehensive analysis of how judges follow diverse criteria and capture criterion-level preference conflicts.
@@ 3.1 Task Formulation
Existing multimodal judge benchmarks [li2025vl, yasunaga2025multimodal, zhang2025mm, chen2024mllm] typically consist of preference pairs $(q,l_{a},l_{b},y)$, where $q$ is a multimodal prompt containing an image and a textual query, $l_{a}$ and $l_{b}$ are candidate responses, and $y\in\{a,b\}$ denotes the golden preference by human annotation or rule-based verification. However, such one-dimensional annotations cannot capture the multifaceted nature of human evaluation—different criteria may lead to conflicting judgments for the same response pair. Multi-Crit therefore extends this formulation to include criterion-specific preferences:
@@ 3.2 Prompt and Response Pair Curation
MMCrit primarily covers two major scopes of multimodal judgment: (1) open-ended content generation, where responses are free-form and traditional fixed metrics are limited; and (2) reasoning tasks, where the judge model evaluates the quality of model-generated reasoning processes leading to objectively verifiable answers.
Prompt Sources.
For open-ended generation, we select images from ImageInWords [garg2024imageinwords] and DOCCI [onoe2024docci] evaluation set for captioning and incorporate image–question pairs from the WildVision-Bench and -Battle datasets [lu2024wildvision], which contain diverse, in-the-wild visual conversation queries. For verifiable reasoning, we draw from reasoning-dense domains, including MathVerse [zhang2024mathverse] (math), MM-K12 [meng2025mm] (STEM), EMMA-mini [hao2025can] (math, science, coding), and VisualPuzzles [song2025visualpuzzles] (visual analogy and puzzle solving).
Response Generation and Pair Curation.
To minimize data contamination, we generate candidate responses using 11 high-performing models, covering both proprietary (e.g. GPT-4o [hurst2024gpt], Gemini-2.5-Flash [comanici2025gemini]) and open-source families (e.g., LLaMA3.2-V [grattafiori2024llama], Qwen2.5-VL [bai2025qwen2], InternVL3 [chen2024internvl]); details in Appendix. We then form two types of response pairs: cross-model pairs, generated by randomly selecting two different LMMs, and intra-model pairs, produced by the same model via temperature sampling. For the latter, we sample $N=5$ responses at temperature 1.0 and select the pair with the largest cosine distance between their MiniLM-L6 [wang2020minilm] embeddings to ensure content diversity. This process yields 3,538 response pairs.
Filtering Challenging Pairs.
We apply a three-stage filtering mechanism to retain challenging yet balanced examples suitable for fine-grained evaluation. (1) Length normalization: We exclude pairs with excessive length disparity, mostly retaining those within [0.7, 1.4] to mitigate superficial bias. (2) Verifiable reasoning filtering: For reasoning tasks, we use GPT-4o-mini to verify both responses and keep only pairs where both are correct or both incorrect, as correctness alone is trivial and our evaluation targets different criteria of reasoning quality. (3) Ensemble-based difficulty filtering: To emphasize nuanced criterion-related differences, we use three strong judge models (GPT-4o, Gemini-2.5-Flash, and Claude-3.7-Sonnet) for initial assessments of overall quality; pairs on which all judges unanimously agree are discarded, while the remaining ones are retained as challenging cases reflecting subtle criterion-level distinctions. A total of 707 response pairs are selected.
@@ 3.3 Criteria-Specific Preference Annotation
Criteria Design.
The criteria in Multi-Crit follow a clear set of principles to ensure balanced coverage of multimodal judging capabilities: (1) Practicality: reflecting common use cases of multimodal judges; (2) Specificity: capturing diverse quality aspects of generated responses while avoiding overlap; and (3) Generality: evaluating fundamental dimensions of model behavior rather than prompt-specific content. Following these principles, we survey existing MLLM-as-a-Judge benchmarks, summarize their adopted criteria, and perform multi-round refinement to derive fivecriteria for open-ended tasks and another five for reasoning tasks, as briefed in Table 2 and detailed in Appendix.
@@
Logic Coherence and Consistency: Follow a clear, step-by-step logic without contradictions or unjustified leaps, and ensures the answer aligns with reasoning.
@@
We recruit nine Ph.D. students majoring in Computer Science as annotators, all with strong backgrounds in multimodal AI and STEM-related domains. Each receives detailed guidelines with clear illustrations of all evaluation criteria. During annotation, annotators are presented with a multimodal instance containing a prompt and two candidate responses. They exclusively evaluate one criterion at a time, determine which response is better (or declare a tie, limited to under 10%), and provide a brief textual justification with supporting evidence. Annotators focus on response elements most relevant to the prompt and indicative of the given criterion. To ensure consistency, we first establish a seed set of 10 open-ended and 10 reasoning examples. Annotators label this set individually, then participate in group discussions led by the project lead to align interpretations and reach consensus. After calibration, each evaluation sample—comprising the prompt, response pair, and set of criteria—is randomly assigned to three annotators for cross-validation in the main phase. The annotation process involved 289 hours of human effort in total. Inter-annotator agreement is quantified using the average pairwise Cohen’s $\kappa$ [artstein2008inter], achieving 0.718 on open-ended generation tasks and 0.805 on verifiable reasoning tasks, indicating substantial and reliable consistency among annotators.
Preference Aggregation and Final Verification.
@@
First, preferences for each criterion are aggregated, retaining only cases where all annotators agree on one model or where two annotators agree and the third declares a tie. Project leads manually review the textual justifications and discard samples with notable inconsistencies or verbose explanations. Subsequently, criterion-level judgments are aggregated at the prompt–response level. Finally, prompt-level filtering primarily retains pairs exhibiting criterion-level conflicts and balances the number of samples across all criteria.
@@
Model Prompt-Level Criterion-Level Pluralisic Conflict Tradeoff Completeness Grounding Hallucination Expressiveness Clarity Avg Open-Source LMMs Qwen2.5-VL-7B-Instruct 9.41 17.28 36.14 56.12 51.70 48.20 64.12 51.82 54.39 LLaVA-OneVision-7B 11.46 12.77 37.86 53.20 48.80 43.10 55.68 52.26 50.61 Llama-3.2-11B-Vision-Instruct 12.71 15.37 41.26 55.75 44.64 48.07 61.14 46.59 51.24 Molmo-7B 17.06 15.97 34.47 48.23 44.20 55.80 58.03 50.57 51.37 Qwen3-VL-8B-Instruct 18.39 18.36 34.47 56.19 47.77 59.12 55.44 52.27 54.16 MiniCPM-V-4.5-8B 18.73 14.77 27.67 53.10 46.43 53.59 65.80 50.00 53.78 Eagle2.5-8B 20.40 23.75 50.97 64.16 54.91 52.49 59.59 54.55 57.14 GLM-4.1V-9B-Thinking 24.08 30.54 55.34 59.73 55.80 61.33 67.36 51.14 59.07 InternVL3.5-8B 25.08 32.34 62.14 61.50 59.38 56.35 69.43 58.52 61.04 InternVL3-8B-Instruct 26.09 30.34 61.17 64.60 62.95 56.35 67.88 56.25 61.61 Qwen2.5-VL-72B-Instruct 28.43 35.53 60.68 69.47 63.39 58.56 70.98 56.82 63.84 MiMo-VL-7B 29.10 39.52 65.53 65.93 62.95 64.09 70.47 53.41 63.37 InternVL3-78B-Instruct 29.10 32.53 56.31 73.01 65.62 56.35 68.91 59.66 64.71 InternVL3.5-38B 30.43 33.73 64.08 71.68 65.18 62.98 63.73 61.93 65.10 Finetuned Judge Models LLaVA-Critic-7B (LlaVA-OV) 11.70 14.17 43.2 56.93 47.32 29.82 50.28 44.08 45.69 R1-Reward-7B (Qwen2.5-VL) 17.73 20.36 45.63 59.29 60.71 49.72 55.44 53.98 55.83 UnifiedReward-7B(Qwen2.5-VL) 18.06 8.38 16.50 57.96 52.23 52.49 57.51 55.68 55.17 LLaVA-Critic-R1-7B(Qwen2.5-VL) 18.39 17.96 39.32 55.31 57.59 46.96 63.73 55.11 55.74 Proprietary LMMs Gemini-2.5-Flash 25.42 31.34 62.14 64.6 60.71 66.30 63.73 57.39 62.55 Gemini-2.5-Pro 28.76 37.92 66.50 65.93 59.82 70.17 62.18 60.23 63.67 GPT-5 29.77 38.52 62.62 69.91 69.64 75.69 58.55 68.75 68.51 o3 31.10 42.71 62.62 66.37 74.55 72.93 63.21 68.75 69.16 GPT-4o 31.44 44.91 66.02 68.14 71.88 65.75 76.17 65.91 69.57 Claude-3.7-Sonnet 31.77 42.32 64.08 71.68 74.55 63.54 66.84 60.23 67.37 o4-mini 32.78 43.11 64.56 76.11 74.55 65.75 69.43 62.50 69.67

### v2

@@ Abstract
{NoHyper}**footnotetext: Core contributors.
@@ 1 Introduction
The recent advancements of large multimodal models (LMMs) [21, 25, 2, 7, 4, 16] in visual understanding [40, 58, 61, 12, 41, 6] and reasoning [28, 64, 18] have underscored the growing need for scalable and reliable multimodal evaluation systems. However, evaluating free-form outputs in open-ended or reasoning-intensive tasks remains inherently challenging, as static or reference-based metrics often fail to generalize and provide limited insight into model behavior. A growing trend has emerged to deploy LMMs themselves as generative judges (LMM-as-a-Judge[65, 44, 53]. Given a multimodal prompt consisting of an image and a user query, model responses, and a set of predefined evaluation criteria, the judge model produces either scalar scores for individual responses or pairwise preference judgments, accompanied by textual justifications for its evaluation. Owing to its scalability and flexibility in defining task-specific criteria, this paradigm has been widely adopted for open-ended evaluation across diverse multimodal benchmarks [27, 12, 64, 57, 60, 21, 18, 37, 51]. A series of works fine-tune open-source models to develop dedicated judge and critic capacities, and have proven effective in providing AI feedback for aligning LMMs through reinforcement fine-tuning [22, 48, 26, 70] or test-time scaling [66, 45, 71].
The reliability of the LMM-as-a-Judge paradigm relies on two key components: (1) consistency with human judgment, and (2) flexibility to follow diverse, task-specific evaluation criteria that capture different aspects of model behavior (e.g., factuality, visual grounding, structural coherence). However, existing multimodal reward and judge benchmarks [23, 68, 3, 63] mainly address the first aspect by providing a single overall preference label for each pair of model responses. This coarse formulation overlooks the evaluation of judge models in producing criterion-specific judgment. Specifically, two responses may exhibit trade-offs across different criteria—for instance, a concise but factually correct answer versus a visually grounded yet slightly hallucinated one. This motivates a key question: Can LMM-based judges follow diverse evaluation criteria and recognize conflicts among criterion-level judgments?
@@
    How to construct evaluation data with multi-criterion human preferences that capture diverse aspects of response quality and reveal intra-sample conflicts? We design a rigorous data curation and preference-annotation pipeline that: (1) collects diverse multimodal prompts spanning both open-ended generation and verifiable reasoning tasks; (2) selects challenging response pairs from various strong LMMs with fine-grained criterion-level quality differences via multi-step filtering; (3) defines clear evaluation criteria covering complementary aspects of multimodal judgment; and (4) gathers high-quality human annotations ensuring reliability and inter-annotator consistency.
@@
Overall, Multi-Crit is the first benchmark for evaluating multimodal judges under pluralistic criteria. Built on a carefully curated dataset and novel metrics, it reveals key limitations and informs more reliable, steerable evaluation.
@@ 2 Related Work
Large Multimodal Model as Judges and Critics. Large multimodal models (LMMs) are widely adopted for evaluating model-generated responses, providing assessments ranging from scalar [21, 61, 64, 12] to pairwise preference [25, 27, 60]. GPT-4V [65] was first shown to align closely with human judges across diverse multimodal evaluation tasks, motivating efforts to develop open-source counterparts. Subsequent works fine-tuned open-source LMMs into either BT-style reward models with dedicated reward heads [37, 62, 67] or generative judges that produce textual justifications and scores [20, 53, 68, 66, 48, 45, 49, 47, 52, 54] in an autoregressive manner. While our benchmark centers on generative judge models—due to their flexible instruction-following abilities inherent in the autoregressive generation process of LMMs—the proposed methodology is broadly applicable across diverse reward modeling paradigms.
Benchmarking Multimodal Judges. Benchmarking reward and judge models is crucial for evaluating the reliability of AI feedback, which underpins both automated evaluation and the enhancement of LMM performance through preference alignment. MLLM-as-a-Judge [3] is the first to evaluate LMM as judges on academic tasks; VL-Rewardbench [23], multimodal-rewardbench [56] and MM-RLHF [68] Benchs extend prompts sources to diverse real-world scenarios, reasoning and safety tasks. Following reward and judge benchmarks in text-only LLMs [19, 38, 72, 69], these multimodal benchmarks evaluate judges by their overall preference between paired responses, measured against human or verifier-provided golden labels. MM-Critic further extends evaluation to assess the correctness of generated textual critiques by leveraging GPT-generated scoring rubrics [63]. However, none of these benchmarks evaluate multimodal judges on their ability to follow diverse criteria for each response pair, and fine-grained, criterion-level evaluation remains underexplored.
Criteria-Following Judgment. When deploying model-based judges [64, 61, 25, 60, 53, 35, 24] for automated evaluation, benchmarks usually define specific criteria such as helpfulness [27], visual grounding [57], or no hallucination [37, 34]. While recent work began evaluating criterion-following of pure-text LLM judges—either by embedding criterion-level differences into responses [33] or summarizing from human justifications [50]—we extend this line of research to the multimodal setting, analyzing how judges follow diverse criteria and capture preference conflicts.
@@ 3.1 Task Formulation
Existing multimodal judge benchmarks [23, 56, 68, 15] typically consist of preference pairs $(q,l_{a},l_{b},y)$, where $q$ is a multimodal prompt containing an image and a textual query, $l_{a}$ and $l_{b}$ are candidate responses, and $y\in\{a,b\}$ denotes the golden preference by human annotation or rule-based verification. However, such one-dimensional annotations cannot capture the multifaceted nature of human evaluation—different criteria may lead to conflicting judgments for the same response pair. Multi-Crit therefore extends this formulation to include criterion-specific preferences:
@@ 3.2 Prompt and Response Pair Curation
Multi-Crit primarily covers two major scopes of multimodal judgment: (1) open-ended generation, where responses are free-form and traditional fixed metrics are limited; and (2) reasoning tasks, where the judge model evaluates the quality of model-generated reasoning processes leading to objectively verifiable answers.
Prompt Sources. For open-ended generation, we select images from ImageInWords [10] and DOCCI [29] evaluation set for captioning and incorporate image–question pairs from the WildVision-Bench and -Battle datasets [27], which contain diverse, in-the-wild visual conversation queries. For verifiable reasoning, we draw from reasoning-dense domains, including MathVerse [64] (math), MM-K12 [28] (STEM), EMMA-mini [13] (math, science, coding), and VisualPuzzles [36] (visual analogy and puzzle solving).
Response Generation and Pair Curation. To minimize data contamination, we generate candidate responses using 11 high-performing models, covering both proprietary (e.g. GPT-4o [16], Gemini-2.5-Flash [8]) and open-source families (e.g., LLaMA3.2-V [11], Qwen2.5-VL [2], InternVL3 [7]); details in Appendix. We then form two types of response pairs: cross-model pairs, generated by randomly selecting two different LMMs, and intra-model pairs, produced by the same model via temperature sampling. For the latter, we sample $N=5$ responses at temperature 1.0 and select the pair with the largest cosine distance between their MiniLM-L6 [42] embeddings to ensure content diversity. This process yields 3,538 response pairs.
Filtering Challenging Pairs. We apply a three-stage filtering mechanism to retain challenging yet balanced examples suitable for fine-grained evaluation. (1) Length normalization: We exclude pairs with excessive length disparity, mostly retaining those within [0.7, 1.4] to mitigate superficial bias. (2) Verifiable reasoning filtering: For reasoning tasks, we use GPT-4o-mini to verify both responses and keep only pairs where both are correct or both incorrect, as correctness alone is trivial and our evaluation targets different criteria of reasoning quality. (3) Ensemble-based difficulty filtering: To emphasize nuanced criterion-related differences, we use three strong judge models (GPT-4o, Gemini-2.5-Flash, and Claude-3.7-Sonnet) for initial assessments of overall quality; pairs on which all judges unanimously agree are discarded, while the remaining ones are retained as challenging cases reflecting subtle criterion-level distinctions. A total of 707 response pairs are selected.
@@ 3.3 Criteria-Specific Preference Annotation
Criteria Design. The criteria in Multi-Crit follow a clear set of principles to ensure balanced coverage of multimodal judging capabilities: (1) Practicality: reflecting common use cases of multimodal judges; (2) Specificity: capturing diverse quality aspects of generated responses while avoiding overlap; and (3) Generality: evaluating fundamental dimensions of model behavior rather than prompt-specific content. Following these principles, we survey existing MLLM-as-a-Judge benchmarks, summarize their adopted criteria, and perform multi-round refinement to derive fivecriteria for open-ended tasks and another five for reasoning tasks, as briefed in Table 2 and detailed in Appendix.
@@
Logic Coherence and Consistency: Follow a clear, step-by-step logic without contradictions or unjustified leaps, and ensure the answer aligns with reasoning.
@@
Rigorous Human Annotation. We recruit nine computer science Ph.D. students as annotators, all with strong backgrounds in multimodal AI and STEM-related domains. Each receives detailed guidelines with clear illustrations of all evaluation criteria. During annotation, annotators are presented with a multimodal instance containing a prompt and two candidate responses. They exclusively evaluate one criterion at a time, determine which response is better (or declare a tie, limited to under 10%), and provide a brief textual justification with supporting evidence. Annotators focus on response elements most relevant to the prompt and indicative of the given criterion. To ensure consistency, we first establish a seed set of 10 open-ended and 10 reasoning examples. Annotators label this set individually, then participate in group discussions led by the project lead to align interpretations and reach consensus. After calibration, each evaluation sample—comprising the prompt, response pair, and set of criteria—is randomly assigned to three annotators for cross-validation in the main phase. The annotation process involved 289 hours of human effort in total. Inter-annotator agreement is quantified using the average pairwise Cohen’s $\kappa$ [1], achieving 0.718 on open-ended generation tasks and 0.805 on verifiable reasoning tasks, indicating substantial and reliable consistency among annotators.
@@
Preference Aggregation and Final Verification. First, preferences for each criterion are aggregated, retaining only cases where all annotators agree on one model or where two annotators agree and the third declares a tie. Project leads manually review the textual justifications and discard samples with notable inconsistencies or verbose explanations. Subsequently, criterion-level judgments are aggregated at the prompt–response level. Finally, prompt-level filtering primarily retains pairs exhibiting criterion-level conflicts and balances the number of samples across all criteria.
@@
Model Prompt-Level Criterion-Level Pluralisic Conflict Tradeoff Completeness Grounding Hallucination Expressiveness Clarity Avg Open-Source LMMs Qwen2.5-VL-7B-Instruct 9.41 17.28 36.14 56.12 51.70 48.20 64.12 51.82 54.39 LLaVA-OneVision-7B 11.46 12.77 37.86 53.20 48.80 43.10 55.68 52.26 50.61 Llama-3.2-11B-Vision-Instruct 12.71 15.37 41.26 55.75 44.64 48.07 61.14 46.59 51.24 Molmo-7B 17.06 15.97 34.47 48.23 44.20 55.80 58.03 50.57 51.37 Qwen3-VL-8B-Instruct 18.39 18.36 34.47 56.19 47.77 59.12 55.44 52.27 54.16 MiniCPM-V-4.5-8B 18.73 14.77 27.67 53.10 46.43 53.59 65.80 50.00 53.78 Eagle2.5-8B 20.40 23.75 50.97 64.16 54.91 52.49 59.59 54.55 57.14 GLM-4.1V-9B-Thinking 24.08 30.54 55.34 59.73 55.80 61.33 67.36 51.14 59.07 InternVL3.5-8B 25.08 32.34 62.14 61.50 59.38 56.35 69.43 58.52 61.04 InternVL3-8B-Instruct 26.09 30.34 61.17 64.60 62.95 56.35 67.88 56.25 61.61 Qwen2.5-VL-72B-Instruct 28.43 35.53 60.68 69.47 63.39 58.56 70.98 56.82 63.84 MiMo-VL-7B 29.10 39.52 65.53 65.93 62.95 64.09 70.47 53.41 63.37 InternVL3-78B-Instruct 29.10 32.53 56.31 73.01 65.62 56.35 68.91 59.66 64.71 InternVL3.5-38B 30.43 33.73 64.08 71.68 65.18 62.98 63.73 61.93 65.10 Finetuned Judge Models LLaVA-Critic-7B (LlaVA-OV) 11.70 14.17 43.20 56.93 47.32 29.82 50.28 44.08 45.69 R1-Reward-7B (Qwen2.5-VL) 17.73 20.36 45.63 59.29 60.71 49.72 55.44 53.98 55.83 UnifiedReward-7B(Qwen2.5-VL) 18.06 8.38 16.50 57.96 52.23 52.49 57.51 55.68 55.17 LLaVA-Critic-R1-7B(Qwen2.5-VL) 18.39 17.96 39.32 55.31 57.59 46.96 63.73 55.11 55.74 Proprietary LMMs Gemini-2.5-Flash 25.42 31.34 62.14 64.60 60.71 66.30 63.73 57.39 62.55 Gemini-2.5-Pro 28.76 37.92 66.50 65.93 59.82 70.17 62.18 60.23 63.67 GPT-5 29.77 38.52 62.62 69.91 69.64 75.69 58.55 68.75 68.51 o3 31.10 42.71 62.62 66.37 74.55 72.93 63.21 68.75 69.16 GPT-4o 31.44 44.91 66.02 68.14 71.88 65.75 76.17 65.91 69.57 Claude-3.7-Sonnet 31.77 42.32 64.08 71.68 74.55 63.54 66.84 60.23 67.37 o4-mini 32.78 43.11 64.56 76.11 74.55 65.75 69.43 62.50 69.67